# supplement_arXiv

### v2

@@ 3.2 Regularized Mahalanobis distance for density data
$M^{2}_{\alpha,k,B^{2}}(X,Y;C)=$ $~{}||X_{\mathrm{st},C}^{{\alpha},k}\ominus Y_{\mathrm{st},C}^{{\alpha},k}||_{B^{2}}^{2},$
@@
The following proposition gives a distribution of the squared RDMD under the assumption of Gaussianity. It also describes how to detect potential outliers by comparing the calculated Mahalanobis distance to an appropriate quantile from the reference distribution.
Proposition 3.4
Let $X\in\mathcal{B}^{2}(I)$ be a constant-mean random density such that $\mathrm{clr}(X)$ is a Gaussian random process. Then, for any $\alpha>0$ and $k\in\mathbf{N}_{0}$
where $\eta_{i}\sim\chi^{2}(1)$, $i\geq 1$ are mutually independent and also independent of $y\sim\chi^{2}(k)$, and $\lambda_{i}$, $i\geq 1$ is the $i$th eigenvalue $C_{B^{2},X}$.
@@ 4.1 Robust Bayes covariance estimation
$\hat{\mu}_{B^{2},H}=\frac{1}{h}\odot\bigoplus_{i\in H}X_{i},\quad\hat{C}_{B^{2},H,c_{H}}=c_{H}\frac{1}{h}\odot\bigoplus_{i\in H}(X_{i}\ominus\hat{\mu}_{B^{2},H})\otimes_{B^{2}}(X_{i}\ominus\hat{\mu}_{B^{2},H}).$
The constant $c_{H}$ is a scaling factor to adjust for trimming and is determined under the assumption of Gaussianity. For that purpose, we present the next corollary which is an analogue of Proposition 3.4 on the sample level.
@@
For more details on the necessity of scaling the trimmed covariance, see, e.g., Boudt et al. (2020); Oguamalam et al. (2024).
Using the notion of RDMD as a centrality measure, we choose the trimming subset $H_{0}$ as a solution to
@@
Oguamalam et al. (2024) (Lemma 3.1) guarantees that the implicit equation (9) has a solution.
@@
$c_{H_{0}}=\mathrm{med}\left\{M_{\alpha,k,B^{2}}^{2}(X_{i},\hat{\mu}_{B^{2},H_{0}};\hat{C}_{B^{2},H_{0},c_{H_{0}}}),\,i=1,\ldots,n\right\}/\mathrm{med}\left\{y+\sum_{i=k+1}^{\infty}\frac{\lambda_{i}^{2}}{(\lambda_{i}+\alpha)^{2}}\eta_{i}\right\},$
where $\eta_{i}\sim\chi^{2}(1)$, $i\geq 1$, are mutually independent and are independent of $y\sim\chi^{2}(k)$ for $k\geq 1$.
@@ 4.2 Algorithm
We solve (9) iteratively, replacing the RDMD in the $j$th iteration with its estimate using the $(j-1)$st optimal subset:
@@
where $\hat{\mu}_{B^{2},H_{j-1}}$ and $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$ are the trimmed sample mean and covariance operator based on the subset $H_{j-1}$, respectively. Thus, in the $j$th step of the procedure, we identify $h$ observations with the smallest regularized Mahalanobis distances to the robust mean estimate $\hat{\mu}_{B^{2},H_{j-1}}$, w.r.t. the current robust covariance estimate $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$.
In practice, it is important to note that we do not observe smooth density functions directly. Instead, we either see discrete realizations of each underlying functional observation or the data sets used to estimate these densities. A common approach for density estimation is kernel-based density estimation, where densities are estimated on an arbitrarily dense, regular grid. Let therefore $X_{i}(t):t\in\{t_{1},\dots,t_{p}\}$, $i=1,\dots,n$, be a discretely observed sample.
Since we observe only a discrete set of functional values for each function $X_{i}$, $i=1,\dots,n$, both $X_{i}$ and $\hat{\mu}_{B^{2},H_{j-1}}$ are represented as large vectors. The covariance $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$ is similarly replaced with a matrix of sample estimates of the pointwise covariances at observed points. If $p$ is chosen large enough, however, functional inner products used in estimating the RDMD can be well approximated by integral sums; see e.g. Ostebee and Zorn (2002) for more detail on integral approximations. Once robust pointwise estimates of the covariance (and PCs) are obtained, interpolation can be applied to extend these estimates over the entire domain, though that lies beyond the scope of this paper.
The strategy for finding the solution to (10), and consequentially robust principal components, results in a concentration-step-type algorithm (Rousseeuw and Driessen, 1999) and is presented in Algorithm 1, where for simplicity of notation we assume $p\geq n$ and $k\geq 1$.
@@
Input : sample $X_{1},\ldots,X_{n}\in\mathcal{B}^{2}$ observed at $p$ time points $t_{1},\dots,t_{p}$, an initial subset $H_{1}\subset\{1,\dots,n\}$, $|H_{1}|=h$, regularization parameter $\alpha>0$, projection dimension $k\leq n$, and the tolerance level $\varepsilon_{k}>0$;
1 $Y_{i}\leftarrow\mathrm{clr}(X_{i})$, $i=1,\dots,n$; do
@@ 2       $H_{0}\leftarrow H_{1}$;
3       Sample mean: $\bar{Y}_{H_{0}}\leftarrow\frac{1}{h}\sum_{i\in H_{0}}Y_{i}$;
4       Non-scaled covariance: $\hat{C}_{\mathrm{clr},H_{0}}\leftarrow\frac{1}{h}\sum_{i\in H_{0}}\left(Y_{i}-\bar{Y}_{H_{0}}\right)\otimes\left(Y_{i}-\bar{Y}_{H_{0}}\right)$;
5       Calculate spectral decomposition $\hat{C}_{\mathrm{clr},H_{0}}=\sum_{i\geq 1}\hat{\lambda}_{i,H_{0}}\hat{\xi}_{i,H_{0}}\otimes\hat{\xi}_{i,H_{0}}$;
@@ 8             $c_{0,H_{0}}\leftarrow c_{1,H_{0}}$;
9             Set $\hat{V}(k)\leftarrow I-\sum_{i=1}^{k}\hat{\xi}_{i,H_{0}}\otimes\hat{\xi}_{i,H_{0}}$;
10             Standardization: $Y_{i,\mathrm{st},H_{0}}^{\alpha,k}=\left(\hat{C}_{\mathrm{clr},H_{0}}+\alpha/c_{0,H_{0}}\hat{V}(k)\right)^{-1}\hat{C}_{\mathrm{clr},H_{0}}^{1/2}(Y_{i}-\bar{Y}_{H_{0}})$;
11             $d_{i,H_{0},c_{0}}^{2}\leftarrow\|Y_{i,\mathrm{st},H_{0}}^{\alpha,k}\|^{2}$;
12             Generate $\eta_{k+1},\dots,\eta_{n}\sim\chi^{2}(1),\quad y\sim\chi^{2}(k)$;
13             $c_{1,H_{0}}\leftarrow\frac{\mathrm{med}\left\{d_{i,H_{0},c_{0}}^{2}:i=1,\dots,n\right\}}{\mathrm{med}\left\{y+\sum_{i=k+1}^{\infty}\frac{\lambda_{i}^{2}}{(\lambda_{i}+\alpha)^{2}}\eta_{i}\right\}}$;
14
@@ 17       Set $H_{1}\leftarrow\{i_{1},\dots,i_{h}\}$;
18
19while $H_{0}\neq H_{1}$;
20Solve symmetric eigenproblem $h^{-1}c_{1,H_{1}}\hat{C}_{\mathrm{clr},H_{1}}(\xi_{j})=\lambda_{j}\xi_{j}$;
21 Set $\hat{\xi}_{j,H_{\mathrm{opt}}}\leftarrow\xi_{j}$, $\hat{\zeta}_{j,H_{\mathrm{opt}}}=\mathrm{clr}^{-1}(\xi_{j})$ and $\hat{\lambda}_{j,H_{\mathrm{opt}}}\leftarrow\lambda_{j}$;
22
Output : Optimal subset $H_{\text{opt}}:=H_{1}$, scaling factor $c_{H_{0}}$, squared robust RDMD $(c_{1,H_{0}}^{-1}d_{(1,H_{0},c_{1})}^{2},\dots c_{1,H_{0}}^{-1}d_{(n,H_{0},c_{1})}^{2})$, robust eigenvalues $\{\hat{\lambda}_{j,H_{\mathrm{opt}}}\}_{j\geq 1}$ and corresponding robust FPCs $\{\hat{\zeta}_{j,H_{\mathrm{opt}}}\}_{j\geq 1},\{\hat{\xi}_{j,H_{\mathrm{opt}}}\}_{j\geq 1}$ in Bayes and clr space, respectively.
Algorithm 1 RDPCA
Alternatively, in the case where the discretely observed data has been converted into a functional form, in particular via finite basis smoothing, an adaptation of Algorithm 1 involves calculations to be based on the matrix of the coefficients of the basis expansion, and can be done as described in Oguamalam et al. (2024).
@@
The regularized Mahalanobis distance depends on the tuning parameters $\alpha>0$ and $k\in\mathbb{N}_{0}$. The choice of $\alpha$ balances the smoothness of the regularized data with how well it fits the observed data. If $\alpha$ is too large, over-regularization will result in poor fit, with large residuals $\|\hat{C}_{B^{2},H_{\mathrm{opt}}}^{1/2}(X_{i,\mathrm{st},\hat{C}_{B^{2},H_{\mathrm{opt}}}}^{{\alpha},k})\ominus X_{i})\|_{B^{2}}^{2}$, $i=1,\ldots,n$. Conversely, if $\alpha$ is too small, the solution will overfit data errors, leading to large values of $\|\hat{W}(k)X_{i,\mathrm{st},\hat{C}_{B^{2},H_{\mathrm{opt}}}}^{{\alpha},k}\|^{2}$.
To determine $\alpha$, we adapt the approach from Oguamalam et al. (2024), iteratively adjusting $\alpha$ so that the Bayes covariance of the standardized data has eigenvalues close to zero (for the latter part of the spectrum) or some fixed constant (for the signal part), thus maximizing the signal-to-noise ratio.
On the other hand, we can think of $k$ as a certain over-smoothing parameter that prevents important signal features from being over-smoothed. Hron et al. (2016) argues how for densities from the $k_{0}$ exponential family (with $k_{0}$ parameters), the number of principal components belonging to strictly positive eigenvalues is precisely $k_{0}$. E.g., for the Gaussian family with the known mean $k_{0}=1$; see Hron et al. (2016) for more details. Therefore, in the Sections 5 and 6 we fix $k=1$.

### v3

@@ 3.2 Regularized Mahalanobis distance for density data
[…]
@@
One can show that the distribution of the squared RDMD under the assumption of Gaussianity is an infinite weighted sum of independent $\chi^{2}(1)$ variables. For the sake of conciseness, the formal result is given in Proposition LABEL:prop:Mah_distribution in Supplement LABEL:supp:aux_results.
@@ 4.1 Robust Bayes covariance estimation
$\hat{\mu}_{B^{2},H}=\frac{1}{h}\odot\bigoplus_{i\in H}X_{i},\quad\hat{C}_{B^{2},H,c_{H}}=c_{H}\frac{1}{h}\odot\bigoplus_{i\in H}(X_{i}\ominus\hat{\mu}_{B^{2},H})\otimes_{B^{2}}(X_{i}\ominus\hat{\mu}_{B^{2},H}).$ (8)
The constant $c_{H}$ is a scaling factor to adjust for trimming and is determined under the assumption of Gaussianity. For that purpose, we use the corollary of Proposition LABEL:prop:Mah_distribution.
@@
For more details on the necessity of scaling the trimmed covariance, see, e.g., Boudt et al. (2020); Oguamalam et al. (2024). Using the notion of RDMD as a centrality measure, we choose the trimming subset $H_{0}$ as a solution to
@@
(Oguamalam et al., 2024, Lemma 3.1) guarantees that the equation (10) has a solution.
@@
$c_{H_{0}}=\mathrm{med}\left\{M_{\alpha,k,B^{2}}^{2}(X_{i},\hat{\mu}_{B^{2},H_{0}};\hat{C}_{B^{2},H_{0},c_{H_{0}}}),\,i=1,\ldots,n\right\}/\mathrm{med}\left\{y+\sum_{i=k+1}^{\infty}\frac{\hat{\lambda}_{i}^{2}}{(\hat{\lambda}_{i}+\alpha)^{2}}\eta_{i}\right\},$
where $\eta_{i}\sim\chi^{2}(1)$, $i\geq 1$, are mutually independent and are independent of $y\sim\chi^{2}(k)$ for $k\geq 1$. We note that, for the purpose of PCA, the scaling is redundant. We include it, however, for completeness.
Remark 4.1
Corollary 4.1 can be used to detect potential outliers by comparing the calculated RDMD to an appropriate quantile from the limiting reference distribution; see Remark 2.1. However, we emphasize how RDMD and the underlying robust estimation do not rely on Gaussianity. In practice, RDMD values can highlight atypical observations, with non-parametric cutoffs, like, e.g., empirical quantiles or boxplots, used without distributional assumptions. The Gaussianity is invoked only to enable an automated, theoretically grounded cutoff for outlier detection, which may be useful in applications such as re-weighting in robust covariance estimation, as e.g. in Rousseeuw and Driessen (1999); Berrendero et al. (2020); Oguamalam et al. (2024).
While general consistency results under a non-contaminated model like those of Croux and Haesbroeck (1999); Butler et al. (1993) for the MCD are beyond the scope of this work, we provide a partial result under a clean model, where the trimming proportion grows sublinearly with the sample size. In this setting, we show that the robust estimators (8) and the associated principal functions converge to the population counterparts. These results are given in Lemma LABEL:sup_lemma:consistency in Supplement Section LABEL:supp:aux_results, along with simulation studies that empirically support this convergence in broader scenarios; see Figures LABEL:fig:additionalResults_consistency_sim1 and LABEL:fig:additionalResults_consistency_sim2.
@@ 4.2 Algorithm
We solve (10) iteratively, replacing the RDMD in the $j$th iteration with its estimate using the (j1)(j-1)st optimal subset:
@@
where $\hat{\mu}_{B^{2},H_{j-1}}$ and $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$ are the trimmed sample mean and covariance operator based on the subset $H_{j-1}$, respectively. Thus, in the $j$th step of the procedure, we identify $h$ observations with the smallest regularized Mahalanobis distances to the robust mean estimate $\hat{\mu}_{B^{2},H_{j-1}}$, w.r.t. the current robust covariance estimate $\hat{C}_{B^{2},H_{j-1},c_{H_{j-1}}}$. However, in practice, we don’t directly observe functions, but rather discrete realizations of the underlying functional observations, or datasets from which such functions are estimated, which is often the case with probability density functions. Let therefore
$X_{i}(t):t\in\{t_{1},\dots,t_{p}\},\,i=1,\dots,n,$ (12)
represent such a discretely observed sample. When the grid is sufficiently dense, the functional inner products required in computing the RDMD can be accurately approximated by integral sums; see, e.g., Ostebee and Zorn (2002). Accordingly, both $X_{i}$ and related quantities such as $\hat{\mu}_{B^{2},H}$ are treated as vectors, and the sample covariance operator $\hat{C}_{B^{2},H,c_{H}}$ is replaced by a matrix of pointwise covariances. The pseudocode for obtaining the robust pointwise estimators of the mean, covariance, and principal functions is given in Algorithm LABEL:alg:alg1 in Supplement Section LABEL:supp:aux_results. After obtaining robust estimates on the observed grid, one may interpolate these estimates to obtain a full representation over the domain, though this step lies outside the scope of this work.
Alternatively, one can first convert discretized observations into a functional representation by projecting them onto a suitable basis. This enables evaluating each function at any point $t\in I$, not just at the sampled time points. Using a fixed number $M\in\mathbb{N}$ of chosen basis functions $\phi_{1}(t),\ldots,\phi_{M}(t):I\rightarrow\mathbb{R}^{+},j=1,\ldots,M,$ each observation can be approximated as:
$X_{i}^{M}(t)=\bigoplus_{j=1}^{M}C_{i,j}\odot\phi_{j}(t),\quad i=1,\dots,n,$ (13)
where $C_{i,j}]\in\mathbb{R}^{n\times M}$ is the matrix of basis coefficients, and the number of basis functions $M$ is typically chosen to strike a balance between accurately capturing the observed data and enforcing smoothness to mitigate noise. While this representation is exact for functions of finite rank, it can also serve as an additional smoothing technique when the basis and the number of components are appropriately chosen. An (orthonormal) basis of our choice is the compositional spline basis introduced in Machalová et al. (2021), designed specifically for the representation of the relative functional data, thus satisfying the unit integral constraint.
Proposition 4.1 shows that the RDMD can be computed exactly in this framework, using only the matrix of basis coefficients $C_{i,j}]$. The corresponding routine for robust covariance and SPCA estimation is given in Algorithm 1.
Proposition 4.1
Let $\boldsymbol{\Phi}(t)=(\phi_{1}(t),\dots,\phi_{M}(t))^{\prime}$, $t\in I$ be an orthonormal basis, and let $X_{i},i=1,\dots,n$ admit a rank-M representation as in (13), where $\textbf{C}=[C_{i,j}]\in\mathbb{R}^{n\times M}$ is the the matrix of basis coefficients. Then, for $H\subseteq\{1,\dots,n\}$, $|H|=h,$ and parameters $\alpha>0$, $k\in\mathbb{N}_{0}$ $\hat{M}^{2}_{\alpha,H,B^{2}}(X_{i},\hat{\mu}_{B^{2},H};\hat{C}_{B^{2},H,c_{H}})=(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H})^{\prime}\textbf{C}\textbf{C}_{H}(\textbf{C}_{H}+\alpha/c_{H}\textbf{W}(k))^{-2}\textbf{C}\,^{\prime}(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H}),$ where $\textbf{C}_{H}=\frac{1}{h}\textbf{C}\,^{\prime}(\textbf{I}_{n,H}-\frac{1}{h}\textbf{J}_{n,H})\textbf{C}\in\mathbb{R}^{M\times M}$, $\mathbf{1}_{H}\in\mathbb{R}^{n}$ is such that $(\mathbf{1}_{H})_{i}=1$ if $i\in H$ and 0 otherwise, $\textbf{J}_{n,H}=\mathbf{1}_{H}\mathbf{1}_{H}^{\prime}$, and $\textbf{I}_{n,H}=\mathrm{diag}(\mathbf{1}_{n,H})\in\mathbb{R}^{n\times n}$ is a diagonal matrix with $\mathbf{1}_{H}$ on its diagonal. $\textbf{W}(k)$ is a projection matrix onto the space spanned by the span of the first $k$ eigenvectors of $\textbf{C}_{H}$; if $k=0$, $\textbf{W}(k)=\textbf{I}_{M}.$
Remark 4.2
Our framework assumes some level of smoothness in the underlying functional data, an assumption embedded via regularization. In practice, this means that the methodology can also be justifiably applied to data observed on a dense, regular grid without requiring basis expansion. In fact, pre-smoothing or excessive preprocessing may mask outlying behavior, potentially interfering with outlier detection (Hubert et al., 2015). Therefore, both discretized and functional representations are supported within our framework—the choice between them should depend on the nature of the data analyst’s preference.
@@
Input : Orthonormal basis $\boldsymbol{\Phi}=(\phi_{1},\dots,\phi_{M})\in\mathcal{B}^{2}$, sample ${X}_{1},\dots,{X}_{n}\in\mathcal{B}^{2}$, subset size $h$, initial subset $H_{1}\subset\{1,\dots,n\}$ with $|H_{1}|=h$, regularization parameter $\alpha>0$, projection dimension $k\leq n$, and tolerance level $\varepsilon_{k}>0$;
1
[…]
3 Express $Y_{i}$ in basis $\tilde{\boldsymbol{\Phi}}$ as $Y_{i}=\sum_{j=1}^{M}C_{i,j}\tilde{\phi}_{j}$, and collect $\boldsymbol{C}=[C_{i,j}]\in\mathbb{R}^{n\times M}$;
4 do
@@ 5 $H_{0}\leftarrow H_{1}$;
6 $\boldsymbol{C}_{H_{0}}=\frac{1}{h}\boldsymbol{C}^{\prime}(\textbf{I}_{n,H_{0}}-\frac{1}{h}\textbf{J}_{n,H_{0}})\boldsymbol{C}=\sum_{i=1}^{M}{\hat{\lambda}}_{i,H_{0}}\textbf{u}_{i}\textbf{u}_{i}^{\prime}$;
@@ 9    $c_{0,H_{0}}\leftarrow c_{1,H_{0}}$;
10      Set $\tilde{V}(k)\leftarrow I-\sum_{j=1}^{k}\hat{\lambda}_{j,H_{0}}\boldsymbol{u}_{j}\boldsymbol{u}_{j}^{\prime}$;
11    $d_{i,H_{0},c_{0}}^{2}\leftarrow(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H_{0}})^{\prime}\boldsymbol{C}\boldsymbol{C}_{H_{0}}(\boldsymbol{C}_{H_{0}}+\alpha/c_{0,H_{0}}\tilde{V}(k))^{-2}\boldsymbol{C}^{\prime}(\textbf{e}_{i}-\frac{1}{h}\mathbf{1}_{H_{0}})$;
12      Generate $\eta_{k+1},\ldots,\eta_{n}\sim\chi^{2}(1)$, $y\sim\chi^{2}(k)$;
13    $c_{1,H_{0}}\leftarrow{\mathrm{med}\left\{d_{i,H_{0},c_{0}}^{2}:i=1,\dots,n\right\}}/{\mathrm{med}\left\{y+\sum_{i=k+1}^{M}\frac{\hat{\lambda}_{i,H_{0}}^{2}}{(\hat{\lambda}_{i,H_{0}}+\alpha/c_{0,H_{0}})^{2}}\eta_{i}\right\}}$;
@@ 16   Set $H_{1}\leftarrow\{i_{1},\dots,i_{h}\}$;
17
18while $H_{0}\neq H_{1}$;
19Solve symmetric eigenproblem $c_{1,H_{1}}\boldsymbol{C}_{H_{1}}(u_{j})=\hat{\lambda}_{j}u_{j};$
20 Recalculate $\xi_{j}=u_{j}^{\prime}\tilde{\boldsymbol{\Phi}}$ and set $\hat{\xi}_{j,H_{\mathrm{opt}}}\leftarrow\xi_{j},\hat{\zeta}_{j,H_{\mathrm{opt}}}\leftarrow\mathrm{clr}^{-1}(\xi_{j})$ and $\hat{\lambda}_{j,H_{\mathrm{opt}}}\leftarrow\hat{\lambda}$;
Output : Optimal subset $H_{\mathrm{opt}}:=H_{1}$, scaling factor $c_{H_{0}}$, squared robust RDMD $(c_{H_{0}}^{-1}d_{1,H_{0},c_{1}}^{2},\dots c_{H_{0}}^{-1}d_{n,H_{0},c_{1}}^{2})$, robust eigenvalues $\{\hat{\lambda}_{j,H_{\mathrm{opt}}}\}_{j\geq 1}$ and corresponding robust FPCs $\{\hat{\zeta}_{j,H_{\mathrm{opt}}}\}_{j\geq 1},\{\hat{\xi}_{j,H_{\mathrm{opt}}}\}_{j\geq 1}$ in Bayes and clr space, respectively.
Algorithm 1 RDPCA for data expressed in a fixed basis
@@
The regularized Mahalanobis distance depends on the tuning parameters $\alpha>0$ and $k\in\mathbb{N}_{0}$. The choice of $\alpha$ balances the smoothness of the regularized data with how well it fits the observed data. If $\alpha$ is too large, over-regularization will result in poor fit, with large residuals $\|\hat{C}_{B^{2},H_{\mathrm{opt}}}^{1/2}(X_{i,\mathrm{st},\hat{C}_{B^{2},H_{\mathrm{opt}}}}^{{\alpha},k})\ominus X_{i})\|_{B^{2}}^{2}$, $i=1,\ldots,n$. Conversely, if $\alpha$ is too small, the solution will overfit data errors, leading to large values of $\|\hat{W}(k)X_{i,\mathrm{st},\hat{C}_{B^{2},H_{\mathrm{opt}}}}^{{\alpha},k}\|^{2}$. To determine $\alpha$, we adapt the approach from Oguamalam et al. (2024), iteratively adjusting $\alpha$ so that the Bayes covariance of the standardized data has eigenvalues close to zero (for the latter part of the spectrum) or some fixed constant (for the signal part), thus maximizing the signal-to-noise ratio.
On the other hand, we can think of $k$ parameter that prevents important signal features from being over-smoothed. Hron et al. (2016) argues how for densities from the $k_{0}$ - exponential family (with $k_{0}$ parameters), the number of principal components belonging to strictly positive eigenvalues is precisely $k_{0}$. E.g., for the Gaussian family with the known mean $k_{0}=1$; see Hron et al. (2016) for more details. Therefore, in Section 5 we fix $k=1$. In more complex situations where the underlying distribution is unknown, selecting $k$ such that the first $k$ principal functions contain at least $90\%$ of variability seems reasonable. This rule of thumb will be applied in Section 6.
Finally, in the supplement, we provide some ideas about the choice of the subset size $h$.